# A Systematic Literature Review of Spatio-Temporal Graph Neural Network Models for Time Series Forecasting and Classification

### v2

@@ 5.4.2 Datasets
African COVID-19 dataset [art_410] https://www.worldometers.info/coronavirus/
Bonn dataset [art_14] http://dx.doi.org/10.1103/PhysRevE.64.061907 (source paper)
Brazilian COVID-19 data [art_52] https://github.com/lcaldeira/GrafoBrasilCovid
CDC US-State [art_414] https://gis.cdc.gov/grasp/fluview/fluportaldashboard.html
Chickenpox [art_410] https://www.kaggle.com/datasets/die9origephit/chickenpox-cases-hungary
@@
EEG DEAP dataset [art_412] http://www.eecs.qmul.ac.uk/mmv/datasets/deap/download.html
eICU [art_189] https://doi.org/10.1038/sdata.2018.178 (source paper)
European COVID-19 dataset [art_410] https://www.worldometers.info/coronavirus/
ILI IDWR Japan-Prefecture [art_414] https://tinyurl.com/y5dt7stm
ILI US-HHS US-Regional [art_414] https://tinyurl.com/y39tog3h
MIMIC-IV [art_189] https://physionet.org/content/mimiciv/2.0/
Nytimes COVID-19 data [art_50], [art_413] https://github.com/nytimes/covid-19-data
OxCGRT dataset [art_416] https://www.bsg.ox.ac.uk/research/covid-19-government-response-tracker
Parkinson’s Progression Markers Initiative (PPMI) [art_92] http://www.ppmi-info.org/
Japan COVID-19 [art_415] https://github.com/swsoyee/2019-ncov-japan
Parkinson Speech data set (PS) [art_92] https://doi.org/10.1109/jbhi.2013.2245674 (source paper)
SJTU Emotion EEG Dataset (SEED) [art_411], [art_412] https://bcmi.sjtu.edu.cn/home/seed/
Spanish COVID-19 data [art_51] https://cnecovid.isciii.es/
@@
Spanish COVID-19 data [art_414] https://dataforgood.fb.com/tools/disease-prevention-maps/
Spikes and slow waves (SSW) dataset [art_14] http://dx.doi.org/10.1109/ICDM50108.2020.00067 (source paper)
@@
US COVID-19  [^2] This dataset includes multiple sources. [art_415] https://github.com/CSSEGISandData/COVID-19
US COVID-19 [art_416] https://covidtracking.com/data/national
US COVID-19 vaccinations [art_50] https://ourworldindata.org/us-states-vaccinations
@@
US influenza [art_48] https://github.com/aI-area/DVGSN/tree/main/dataset
@@
Only two pairs of papers in this thematic group use the same dataset, but none of the proposed models yield comparable results. Most of the datasets are related to COVID-19 and have a daily time granularity. In addition, some papers (Refs. [art_49, art_52, art_92, art_189, art_413]) include exogenous variables that can be either context-related (e.g., geography, population, health conditions, vaccinations, economic factors), or specific to a single patient (including personal information such as age, gender, and medical conditions). A paper that undertakes an in-depth study of exogenous variables is Ref. [art_416]. The authors categorize these variables as holiday variables, time categorical variables (e.g., categorical week and month, and the number of days and weeks since the start of the dataset), and policy variables (e.g., school and workplace closing, movement restrictions, face coverings, and vaccination policies). They carry on a significance analysis of these variables, and explore the correlation among the significant ones. The COVID-19 policy dataset (OxCGRT) which includes three years of 23 policy indicators across more than 180 countries is available at https://www.bsg.ox.ac.uk/research/covid-19-government-response-tracker.
@@ 5.4.3 Types of models
As for the taxonomy of the proposed models, 11 out of 15 papers use a convolutional GNN architecture. Three papers use a pure attentional model, while another paper presents a hybrid convolutional-attentional architecture. Different approaches are used to define the graph structure. Some papers base their graphs on the geographical distribution of the data (as in Refs. [art_49, art_52, art_92]), and others use models that learn the graph structure directly, as in Refs. [art_48, art_50, art_415, art_416]. There is also a paper that uses a variant of the visibility algorithm [art_14]. In half of the cases, the graph is static, while the dynamic graph approach appears more frequently in the papers from 2024.
In terms of programming languages, the community of the “Health” thematic group uses only Python with PyTorch. As for the coice of the loss functions, MSE is typically used for prediction tasks, and cross entropy is used for classification tasks. Tab. LABEL:t:link_codici_health shows the links to the related source codes.
@@
Backbone-based Dynamic Spatio-Temporal Graph Neural Network (BDSTGNN) [art_415] https://github.com/JkMaoya/BDSTGNN
Dynamic virtual graph significance networks (DVGSN) [art_48] https://github.com/aI-area/DVGSN
Deep learning of contagion dynamics on complex networks [art_51] https://doi.org/10.5281/zenodo.4974521
Ordinary Differential Equation in Graph Neural Networks (GNNODENet) [art_413] https://github.com/xiongzhangxyq/GNNODENet
Recurrent Graph Gate Fusion Transformers (ReGraFT) [art_416] https://github.com/mfriendly/ReGraFT
Sparse spectra graph convolutional network (SSGCNet) [art_14] https://github.com/anonymous2020-source-code/WNFG-SSGCNet-ADMM
Spatial-Temporal Graph Convolutional Network (STGCN) [art_52] https://github.com/lcaldeira/CovidPandemicForecasting
Time-aware Context-Gated Graph Attention Network (T-ContextGGAN) [art_189] https://github.com/OwlCitizen/TContext-GGAN
@@ 5.4.4 Benchmark models
Mathematical and statistical methods Autoregressive integrated moving average (ARIMA) [art_48], [art_49], [art_52], [art_410], [art_414], [art_415]
SIR model [modello_SIR] [art_49], [art_413], [art_415]
Non-GNN machine learning methods Feed-forward neural network (FNN) [art_14], [art_48]
Bi-Hemisphere Domain Adversarial Neural Network (Bi-DANN) [Bi-DANN] [art_411], [art_412]
A Novel Bi-Hemispheric Discrepancy Model for EEG Emotion Recognition (BiHDM) [BiHDM] [art_411], [art_412]
Gated recurrent unit (GRU) [art_410], [art_413], [art_415], [art_416]
Long-short term memory network (LSTM) [art_49], [art_50], [art_410], [art_413]
Recurrent neural network (RNN) [art_189], [art_413], [art_414]
Support vector machine (SVM) [art_189], [art_411]
GNN methods Cross-location attention based graph neural network (ColaGNN) [ColaGNN] [art_414], [art_415], [art_416]
Diffusion convolutional recurrent neural network (DCRNN) [DCRNN] [art_49], [art_413], [art_414], [art_416]
Dynamical graph convolutional neural networks (DGCNN) [DGCNN] [art_411], [art_412]
Graph attention network (GAT) [GAT] [art_48], [art_413]
Graph convolutional network (GCN) [art_49], [art_413]
Regularized Graph Neural Networks (RGNN) [RGNN] [art_411], [art_412]
Spatio-temporal attention network for pandemic prediction (STAN) [STAN] [art_413], [art_415]
Spatial-temporal graph ODE networks (STGODE) [STGODE] [art_413], [art_415]
@@ 5.5 Mobility
The thematic group that includes the largest number of papers identified by our query (66 out of 281) is “Mobility”, which regards the movement of people. It includes urban traffic, air travel, and bicycle demand, among other applications. Given the large number of articles in this group, a more detailed comparison of benchmarks, models, and results can be provided.
@@ 5.5.1 Overview
Early traffic prediction is essential for improving the efficiency of transportation systems, helping drivers plan their trips more effectively, and, in the context of cities, preventing urban congestion. The advent of smart cities infrastructures and transportation systems has contributed to the collection of rich data from road sensors. Making accurate traffic forecasts is challenging, due to constantly changing traffic patterns over time and space, as well as the influence of external factors such as weather conditions and special events. Unlike statistical and non-GNN machine learning models, GNNs can be particularly effective for traffic forecasting due to their ability to model complex relationships in spatial and temporal domain. The increasing number of papers published each year in this thematic group reflects both the growing research interest in this field and the growing confidence in GNN models. The most commonly investigated areas regard urban traffic (approximately 70% of the papers), especially urban traffic flow and urban traffic speed. Among the journals where the selected papers were published the most common are Applied Intelligence by Springer, ACM Transactions on Knowledge Discovery from Data, both by the Association for Computing Machinery, and IEEE Transactions on Intelligent Transportation Systems by the Institute of Electrical and Electronics Engineers, with 6 out of 66 papers each. Most of the corresponding authors are affiliated with Chinese institutions.
@@ 5.5.2 Datasets
Bureau of Transportation Statistics (BTS) [art_434], [art_443], [art_448] https://www.transtats.bts.gov/
CAAC flights [art_129] http://www.caac.gov.cn/index.html
@@
Capital Bike Share [art_446] https://ride.capitalbikeshare.com/system-data
CD-HW [art_126] https://doi.org/10.1109/ACCESS.2020.3027375 (source paper)
DiDi Beijing [art_102], [art_447] https://github.com/didi/TrafficIndex
DiDi Chengdu [art_93], [art_98] https://gaia.didichuxing.com/
@@
Divvy Bike Chicago [art_446] https://ride.divvybikes.com/system-data
Huai’an Bus Transaction [art_452] https://www.pandabus.cn/panda/panda-minibus.html
Los-loop [art_103] https://github.com/lehaifeng/T-GCN/tree/master/data
@@
Lyft bike dataset [art_446] https://www.lyft.com/bikes/bay-wheels/system-data
METR-LA [art_93], [art_99], [art_100], [art_109], [art_110], [art_113], [art_116], [art_120], [art_122], [art_126], [art_128], [art_408], [art_427], [art_430], [art_433], [art_435], [art_438], [art_441], [art_442], [art_444] https://paperswithcode.com/dataset/metr-la
Milan network [art_123] https://github.com/arunasubbiah/milan-telecom-data-modeling
NE-BJ [art_100] https://github.com/tsinghua-fib-lab/Traffic-Benchmark
NYC-Bike [art_94], [art_96], [art_114], [art_446], [art_449], [art_451] https://citibikenyc.com/system-data
NYC SafeGraph [art_449] https://www.safegraph.com/academics
@@
NYC-Taxi [art_96], [art_449] https://www1.nyc.gov/site/tlc/about/tlc-trip-record-data.page
PeMS03 [art_108], [art_113], [art_115], [art_117], [art_125], [art_130], [art_419], [art_435], [art_445], [art_453] https://www.kaggle.com/datasets/elmahy/pems-dataset
PeMS04 (PeMSD4) [art_97], [art_99], [art_104], [art_108], [art_113], [art_115], [art_117], [art_122], [art_125], [3], [art_130], [art_131], [art_132], [art_406], [art_408], [art_419], [art_424], [art_433], [art_435], [art_441], [art_445], [art_448], [art_449], [art_452], [art_453] https://paperswithcode.com/dataset/pems04
PeMS07 (PeMSD7) [art_108], [art_110], [art_113], [art_117], [art_125], [3], [art_130], [art_406], [art_422], [art_424], [art_427], [art_432], [art_435], [art_438], [art_441], [art_444], [art_445], [art_453] https://paperswithcode.com/dataset/pems07
PeMS08 (PeMSD8) [art_97], [art_99], [art_108], [art_115], [art_117], [art_122], [art_125], [art_130], [art_131], [art_132], [art_406], [art_408], [art_424], [art_432], [art_433], [art_435], [art_441], [art_444], [art_445], [art_448], [art_449], [art_452], [art_453] https://paperswithcode.com/dataset/pems08
PEMS-BAY [art_100], [art_109], [art_110], [art_116], [art_120], [art_122], [art_406], [art_408], [art_419], [art_420], [art_422], [art_430], [art_433], [art_435], [art_441], [art_442], [art_444] https://paperswithcode.com/dataset/pems-bay
Q-Traffic [art_124] https://github.com/JingqingZ/BaiduTraffic#Dataset
Seattle loop [art_441] https://github.com/zhiyongc/Seattle-Loop-Data
UVDS [art_128] https://doi.org/10.1007/978-3-030-73280-6_6 (source paper)
@@
TaxiBJ [art_126], [art_451] https://www.microsoft.com/en-us/research/publication/forecasting-citywide-crowd-flows-based-big-data/
@@
Taxi Shenzhen [art_103], [art_135] https://github.com/lehaifeng/T-GCN/tree/master/data
@@
TLC Taxi [art_119] https://www.nyc.gov/site/tlc/about/tlc-trip-record-data.page
Urban-core [art_441] https://topis.seoul.go.kr/
Xiecheng delay dataset [art_443] https://pan.baidu.com/s/1dEPyMGh#list/path\$=\$%2F
@@
The most popular datasets used in this group are those provided by the California Transportation Agency Performance Measurement System (PeMS, available at http://pems.dot.ca.gov/), which include various versions differing in time range, spatial coverage, and number of sensors, and the one called METR-LA, collected from loop detectors along the Los Angeles County highway. The number of sensors is in the hundreds, and such is also the number of nodes in the graph. For a more detailed overview of the benchmark datasets in the traffic domain, see Ref. [review_traffico_con_dataset]. Other common public traffic datasets can be found at https://paperswithcode.com/task/traffic-prediction. As to external data, the most commonly used are data on weather conditions, time information (time of day and day of week) and events (e.g., holidays), as well as information about the road network, like points of interest.
@@
The majority of the papers utilize data with a 5-minute granularity, and provide forecasts for horizons ranging from 5 minutes to 1 hour ahead. In addition to imputation of missing values, in almost half of the cases the pre-processing of the data includes data normalization (min-max or Z-score normalization) and the selection of time windows regarding recent, daily and weekly data (as in 10 out of 66 papers).

### v4

@@ 5.4.2 Datasets
African COVID-19 dataset [270] https://www.worldometers.info/coronavirus/
Bonn dataset [383] http://dx.doi.org/10.1103/PhysRevE.64.061907 (source paper)
Brazilian COVID-19 data [298] https://github.com/lcaldeira/GrafoBrasilCovid
CDC US-State [528], [68] https://gis.cdc.gov/grasp/fluview/fluportaldashboard.html
Chickenpox [270], [160] https://www.kaggle.com/datasets/die9origephit/chickenpox-cases-hungary
CSSE COVID-19 [334], [282], [514] https://github.com/CSSEGISandData/COVID-19
@@
EEG DEAP dataset [119] http://www.eecs.qmul.ac.uk/mmv/datasets/deap/download.html
eICU [434] https://doi.org/10.1038/sdata.2018.178 (source paper)
European COVID-19 dataset [270] https://www.worldometers.info/coronavirus/
ILI IDWR Japan-Prefecture [528], [68] https://tinyurl.com/y5dt7stm
ILI US-HHS US-Regional [528], [68] https://tinyurl.com/y39tog3h
ISRUC-Sleep Subgroup III (ISRUC-S3) [384] https://sleeptight.isr.uc.pt/
Japan COVID-19 [285] https://github.com/swsoyee/2019-ncov-japan
Montreal archive of sleep studies (MASS-SS3) [384] http://www.ceams-carsm.ca/en/MASS
@@
Meta COVID-19 data [528], [160] https://dataforgood.fb.com/tools/disease-prevention-maps/
MIMIC-IV [434] https://physionet.org/content/mimiciv/2.0/
Nytimes COVID-19 data [235], [445] https://github.com/nytimes/covid-19-data
OxCGRT dataset [187] https://www.bsg.ox.ac.uk/research/covid-19-government-response-tracker
Parkinson’s progression markers initiative (PPMI) [498] http://www.ppmi-info.org/
Parkinson speech dataset (PS) [498] https://doi.org/10.1109/jbhi.2013.2245674 (source paper)
SJTU emotion EEG dataset (SEED) [527], [119], [384] https://bcmi.sjtu.edu.cn/home/seed/
Spanish COVID-19 data [290] https://cnecovid.isciii.es/
Spikes and slow waves (SSW) [383] http://dx.doi.org/10.1109/ICDM50108.2020.00067 (source paper)
TUH EEG seizure corpus [23] https://isip.piconepress.com/projects/nedc/html/tuh_eeg/
@@
US COVID-19 Johns Hopkins This dataset includes multiple sources. [285] https://github.com/CSSEGISandData/COVID-19
US COVID-19 tracking project [187] https://covidtracking.com/data/national
US COVID-19 vaccinations [235] https://ourworldindata.org/us-states-vaccinations
@@
US influenza [483] https://github.com/aI-area/DVGSN/tree/main/dataset
@@
Only few subsets of papers in this thematic group use the same dataset, but none of the proposed models yield comparable results. Most of the datasets are related to COVID-19 and have a daily time granularity. In addition, some papers (Refs. [462, 298, 498, 434, 445, 334, 282]) include exogenous variables that can be either context-related (e.g., geography, population, health conditions, vaccinations, economic factors, air quality), or specific to a single patient (including personal information such as age, gender, and medical conditions). A paper that undertakes an in-depth study of exogenous variables is Ref. [187]. The authors categorize these variables as holiday variables, time categorical variables (e.g., categorical week and month, and the number of days and weeks since the start of the dataset), and policy variables (e.g., school and workplace closing, movement restrictions, face coverings, and vaccination policies). They carry on a significance analysis of these variables, and explore the correlation among the significant ones. The COVID-19 policy dataset (OxCGRT) which includes three years of 23 policy indicators across more than 180 countries is available at https://www.bsg.ox.ac.uk/research/covid-19-government-response-tracker.
@@ 5.4.3 Types of models
As for the taxonomy of the proposed models, 18 out of 23 papers use a convolutional GNN architecture. Three papers use a pure attentional model, while two other papers present a hybrid convolutional-attentional architecture. Different approaches are used to define the graph structure. Some papers base their graphs on the geographical distribution of the data (as in Refs. [462, 298, 498, 334, 160]), and others use models that learn the graph structure directly, as in Refs. [483, 235, 285, 187, 282]. There is also a paper that uses a variant of the visibility algorithm [383]. In half of the cases, the graph is static, while the dynamic graph approach appears more frequently in the papers from 2024.
In terms of programming languages, the community of the “Health” thematic group uses only Python with PyTorch. As for the choice of the loss functions, MSE is typically used for prediction tasks, and cross entropy is used for classification tasks. Tab. LABEL:t:link_codici_health shows the links to the related source codes.
@@
Backbone-based dynamic spatio-temporal graph neural network (BDSTGNN) [285] https://github.com/JkMaoya/BDSTGNN
DCRNN for COVID-19 forecasting [334] https://github.com/Covid-19-papers/Recurrent-Graph-Neural-Network-on-COVID-19-and-NASA-Satellite-data
Deep learning of contagion dynamics on complex networks [290] https://doi.org/10.5281/zenodo.4974521
Dynamic virtual graph significance networks (DVGSN) [483] https://github.com/aI-area/DVGSN
Ordinary differential equation in graph neural networks (GNNODENet) [445] https://github.com/xiongzhangxyq/GNNODENet
Hierarchical spatial-temporal framework for COVID-19 trend forecasting (HierST) [514] https://github.com/shun-zheng/HierST
Recurrent graph gate fusion transformers (ReGraFT) [187] https://github.com/mfriendly/ReGraFT
Sparse spectra graph convolutional network (SSGCNet) [383] https://github.com/anonymous2020-source-code/WNFG-SSGCNet-ADMM
Spatial-temporal graph convolutional network (STGCN) [298] https://github.com/lcaldeira/CovidPandemicForecasting
Time-aware context-gated graph attention network (T-ContextGGAN) [434] https://github.com/OwlCitizen/TContext-GGAN
Temporal multiresolution graph neural networks (TMGNN) [160] https://github.com/bachnguyenTE/temporal-mgn
@@ 5.4.4 Benchmark models
Mathematical and statistical methods Autoregressive integrated moving average (ARIMA) [483], [462], [298], [270], [528], [285], [160], [68], [282]
Historical average (HA) [160], [282]
Naive [298], [160]
SIR model [180] [462], [445], [285]
Non-GNN machine learning methods Feed-forward neural network (FNN) [383], [483]
Bi-hemisphere domain adversarial neural network (Bi-DANN) [238] [527], [119]
A novel bi-hemispheric discrepancy model for EEG emotion recognition (BiHDM) [234] [527], [119]
Convolutional neural network and recurrent neural network with residual module (CNNRNN-Res) [421] [528], [68]
Gated recurrent unit (GRU) [270], [445], [285], [187]
Long-short term memory network (LSTM) [462], [235], [270], [445], [38], [384], [160], [68], [282]
Long-short term time series network (LSTNet) [202] [528], [68]
MiniRocket [65] [38], [23]
Random forest (RF) [483], [384]
Recurrent neural network (RNN) [434], [445], [528], [68]
Recurrent neural network with attention (RNN+Att) [528], [68]
Support vector machine (SVM) [434], [527], [384]
Time-Series representation learning framework via Temporal and Contextual Contrasting (TS-TCC) [77] [38], [23]
Time series transformer (TST) [471] [235], [23]
GNN methods Cross-location attention based graph neural network (ColaGNN) [68] [528], [285], [187]
CovidGNN [179] [285], [282]
Diffusion convolutional recurrent neural network (DCRNN) [237] [462], [445], [528], [187], [68]
Dynamical graph convolutional neural networks (DGCNN) [354] [527], [119], [384]
Graph attention network (GAT) [376] [483], [445]
Graph convolutional network (GCN) [462], [445]
Regularized graph neural networks (RGNN) [517] [527], [119]
Spatio-temporal attention network for pandemic prediction (STAN) [94] [445], [285], [282]
Spatio-temporal graph convolutional network (STGCN) [457] [445], [68]
Spatial-temporal graph ODE networks (STGODE) [84] [445], [285]
@@ 5.5 Mobility
The thematic group that includes the largest number of papers identified by our query (102 out of 366) is “Mobility”, which regards the movement of people. It includes urban traffic, air travel, and bicycle demand, among other applications. Given the large number of articles in this group, a more detailed comparison of benchmarks, models, and results can be provided.
@@ 5.5.1 Overview
Early traffic prediction is essential for improving the efficiency of transportation systems, helping drivers plan their trips more effectively, and, in the context of cities, preventing urban congestion. The advent of smart cities infrastructures and transportation systems has contributed to the collection of rich data from road sensors. Making accurate traffic forecasts is challenging, due to constantly changing traffic patterns over time and space, as well as the influence of external factors such as weather conditions and special events. Unlike statistical and non-GNN machine learning models, GNNs can be particularly effective for traffic forecasting due to their ability to model complex relationships in spatial and temporal domain. The increasing number of papers published each year in this thematic group reflects both the growing research interest in this field and the growing confidence in GNN models. The most commonly investigated areas regard urban traffic (approximately $1/4$ of the papers), especially urban traffic flow and urban traffic speed. Among the journals where the selected papers were published the most common are Applied Intelligence by Springer, ACM Transactions on Knowledge Discovery from Data, both by the Association for Computing Machinery, and IEEE Transactions on Intelligent Transportation Systems by the Institute of Electrical and Electronics Engineers, with 6 out of 66 papers each. The three conference venues with the highest number of papers are CIKM with 13 papers, SIGSPATIAL with 9 papers, and KDD with 8 papers.
@@ 5.5.2 Datasets
Baoding dataset [493] https://github.com/usail-hkust/ASeer
BRN+ [63] https://www.microsoft.com/en-us/research/publication/t-drive-trajectory-data-sample/
Bruxelles transport data [320] https://www.kaggle.com/datasets/giobbu/belgium-obu
Bureau of transportation statistics (BTS) [420], [419] https://www.transtats.bts.gov/
CAAC flights [25] http://www.caac.gov.cn/index.html
@@
Capital bike share [89] https://ride.capitalbikeshare.com/system-data
CarPark1918 [171] https://github.com/JIANGYUE61610306/SAGDFN/tree/main/data/model
CD-HW [39] https://doi.org/10.1109/ACCESS.2020.3027375 (source paper)
DiDi Beijing [402], [126] https://github.com/didi/TrafficIndex
DiDi Chengdu [274], [220], [276], [144], [219], [275] https://gaia.didichuxing.com/
Didi Shenzhen [276], [144] https://outreach.didichuxing.com/app-outreach/CTIE
@@
Divvy bike Chicago [89], [135] https://ride.divvybikes.com/system-data
Highways England [127] http://tris.highwaysengland.co.uk/detail/trafficflowdata
Los-loop [263] https://github.com/lehaifeng/T-GCN/tree/master/data
@@
Lyft bike dataset [89] https://www.lyft.com/bikes/bay-wheels/system-data
METR-LA [274], [218], [212], [217], [200], [198], [120], [490], [181], [39], [22], [32], [13], [199], [242], [510], [406], [348], [378], [208], [341], [171], [276], [393], [178], [342], [524], [144], [85], [213], [523], [222], [33], [281], [308], [411] https://www.kaggle.com/datasets/annnnguyen/metr-la-dataset
NE-BJ [212] https://github.com/tsinghua-fib-lab/Traffic-Benchmark
NYC-Bike [428], [211], [245], [89], [388], [259], [95], [422] https://citibikenyc.com/system-data
@@
NYC-Taxi [211], [259], [95], [422], [523], [496] https://www1.nyc.gov/site/tlc/about/tlc-trip-record-data.page
PeMS03 [173], [198], [438], [111], [511], [112], [163], [510], [250], [82], [81], [203], [84], [259], [80], [95], [213], [63], [461], [145] https://www.kaggle.com/datasets/elmahy/pems-dataset
PeMS04 (PeMSD4) [62], [218], [489], [173], [198], [438], [111], [181], [511], [112], [69], [27], [227], [32], [163], [407], [242], [510], [348], [250], [82], [341], [127], [81], [342], [203], [84], [8], [80], [95], [530], [85], [213], [63], [370], [461], [145] https://www.kaggle.com/datasets/elmahy/pems-dataset
PeMS07 (PeMSD7) [173], [200], [198], [111], [511], [112], [227], [466], [407], [13], [352], [510], [406], [348], [208], [250], [82], [178], [81], [203], [84], [80], [95], [85], [213], [370], [461], [145] https://www.kaggle.com/datasets/elmahy/pems-dataset
PeMS08 (PeMSD8) [62], [218], [173], [438], [111], [181], [511], [112], [69], [27], [227], [32], [407], [352], [242], [510], [348], [208], [250], [82], [127], [81], [342], [203], [84], [259], [8], [80], [95], [530], [213], [63], [370], [219], [461], [281], [145] https://www.kaggle.com/datasets/elmahy/pems-dataset
PEMS-BAY [212], [217], [200], [120], [490], [181], [227], [32], [163], [159], [466], [199], [242], [510], [348], [378], [208], [341], [276], [393], [178], [342], [524], [144], [85], [213], [222], [281], [308], [411] https://zenodo.org/records/4263971
Q-Traffic [278] https://github.com/JingqingZ/BaiduTraffic#Dataset
SafeGraph [123] https://www.safegraph.com/
Seattle loop [348], [33] https://github.com/zhiyongc/Seattle-Loop-Data
UVDS [22] https://doi.org/10.1007/978-3-030-73280-6_6 (source paper)
@@
TaxiBJ [39], [388], [49] https://www.microsoft.com/en-us/research/publication/forecasting-citywide-crowd-flows-based-big-data/
@@
Taxi Shenzhen [263], [169] https://github.com/lehaifeng/T-GCN/tree/master/data
@@
TLC Taxi [251] https://www.nyc.gov/site/tlc/about/tlc-trip-record-data.page
Urban-core [348] https://topis.seoul.go.kr/
Xiecheng delay [419] https://pan.baidu.com/s/1dEPyMGh#list/path\$=\$%2F
Zhuzhou dataset [493] https://github.com/usail-hkust/ASeer
@@
The most popular datasets used in this group are those provided by the California Transportation Agency Performance Measurement System (PeMS, available at http://pems.dot.ca.gov/), which include various versions differing in time range, spatial coverage, and number of sensors, and the one called METR-LA, collected from loop detectors along the Los Angeles County highway. The number of sensors is in the hundreds, and such is also the number of nodes in the graph. For a more detailed overview of the benchmark datasets in the traffic domain, see Ref. [167]. Other common public traffic datasets can be found at https://web.archive.org/web/20241212092408/https://paperswithcode.com/task/traffic-prediction. As to external data, the most commonly used are data on weather conditions, time information (time of day and day of week) and events (e.g., holidays), as well as information about the road network, like points of interest.
@@
The majority of the papers utilize data with a 5-minute granularity, and provide forecasts for horizons ranging from 5 minutes to 1 hour ahead. In addition to imputation of missing values, in almost half of the cases the pre-processing of the data includes data normalization (min-max or Z-score normalization) and the selection of time windows regarding recent, daily and weekly data.